# Active Learning for Continual Learning: Keeping the Past Alive in the Present

### v1

@@ Abstract
Continual learning (CL) enables deep neural networks to adapt to ever-changing data distributions. In practice, there may be scenarios where annotation is costly, leading to active continual learning (ACL), which performs active learning (AL) for the CL scenarios when reducing the labeling cost by selecting the most informative subset is preferable. However, conventional AL strategies are not suitable for ACL, as they focus solely on learning the new knowledge, leading to catastrophic forgetting of previously learned tasks. Therefore, ACL requires a new AL strategy that can balance the prevention of catastrophic forgetting and the ability to quickly learn new tasks. In this paper, we propose AccuACL, Accumulated informativeness-based Active Continual Learning, by the novel use of the Fisher information matrix as a criterion for sample selection, derived from a theoretical analysis of the Fisher-optimality preservation properties within the framework of ACL, while also addressing the scalability issue of Fisher information-based AL. Extensive experiments demonstrate that AccuACL significantly outperforms AL baselines across various CL algorithms, increasing the average accuracy and forgetting by $23.8\%$ and $17.0\%$, respectively, in average.
@@ 1 Introduction
(a) Accumulative informativeness in ACL.             (b) AL performance on CL benchmarks.
Figure 1: Overview of AccuACL; (a) Unlike conventional AL strategies that only focus on the new task and cause catastrophic forgetting, AccuACL balances the prevention of catastrophic forgetting and the ability to learn new tasks quickly, by defining the accumulative informativeness; (b) shows the catastrophic forgetting of the conventional AL strategies on SplitCIFAR100.
@@
To answer the aforementioned question, we formulate the accumulative informativeness as a novel standard for informativeness in ACL, which balances the prevention of catastrophic forgetting and the ability to quickly learn new tasks. It enables assessing an example’s informativeness with respect to both past and new tasks. Then, we introduce AccuACL (Accumulated informativeness-based Active Continual Learning), an algorithm based on the theoretical analysis of the Fisher information-based AL (Sourati et al., 2017, Zhang and Oles, 2000), that maximizes the accumulated informativeness. We model the accumulated informativeness via the Fisher information matrix, through the approximation with a small memory buffer commonly adopted by many rehearsal-based CL methods, the model parameters, and the unlabeled data pool for the new task, as illustrated in Figure 1(a). As a result, AccuACL becomes to prefer the examples that comprehensively contain the four important features (i.e., Features 1$\sim$4) in Figure 1(a). Furthermore, because Fisher information-based AL is a combinatorial optimization problem, we develop a query algorithm based on two theoretical properties that an optimal labeled subset should possess. To the best of our knowledge, this is the first ACL study that offers the ability to avoid catastrophic forgetting.
Extensive experiments with four different CL methods on three CL benchmarks, SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet, show that AccuACL significantly boosts CL approaches, outperforming conventional AL baselines by $23.8\%$ and $17.0\%$, in terms of average accuracy and forgetting, respectively. As shown in Figure 1(b), AccuACL always achieves dominance in both metrics throughout the entire sequence of tasks. This overall superior performance is attributed to the higher performance especially on the examples for past tasks. The source code is available at https://github.com/kaist-dmlab/AccuACL.
@@ 2 Related Work
Active Learning (AL). AL is a research field that focuses on the selection of unlabeled data points for labeling by an oracle, which provides supervision, especially in domains where labeling requires significant costs Ren et al. (2021), Tharwat and Schenck (2023). This process allows us to optimize model performance under labeling budget constraints. Many strategies exist for selecting informative examples from unlabeled data. Uncertainty-based approaches measure model prediction uncertainty and select the most uncertain examples (Roth and Small, 2006, Settles, 1995, Wang and Shang, 2014). Diversity-based methods prioritize diverse examples to reduce redundancy between selected examples and capture a broader range of patterns and complexities in the data pool (Sener and Savarese, 2017). Hybrid approaches use gradient space embedding to select diverse and uncertain examples (Ash et al., 2019). Fisher information-based methods measure the asymptotic value of unlabeled data with theoretical guarantee (Sourati et al., 2017, Zhang and Oles, 2000). Notably, Kirsch and Gal (2022) demonstrate that modern AL algorithms optimize the same Fisher-based objective.
@@
Continual Learning (CL). CL has gained interest as a way to adapt models to new tasks over time. Numerous research has investigated different methods to address the stability-plasticity dilemma (Mermillod et al., 2013). Rehearsal-based methods store (Aljundi et al., 2019, Buzzega et al., 2020, Caccia et al., 2022, Rahaf and Lucas, 2019, Rolnick et al., 2019, Liang and Li, 2024, Kim et al., 2023a, Lin et al., 2024) or generate (Shin et al., 2017) a subset of examples from past tasks at a constant cost. These examples are then replayed for the new task to retain past knowledge. Regularization-based methods discourage significant changes to model parameters that are essential for past tasks (Aljundi et al., 2018, Kirkpatrick et al., 2017). Dynamic-structure-based methods generate distinct modules to augment the ability to learn new tasks (Yan et al., 2021, Zhou et al., 2023, Wang et al., 2023). Rehearsal-free or prompt-based methods are gaining popularity in the CL field, using pre-trained models and prompt tuning to adjust to data distribution shifts (Wang et al., 2022b; c).
@@
Active Continual Learning (ACL). There has been a lack of research effort on ACL. Vu et al. (2023) support our goal of tackling the impracticality of fully-supervised datasets in real-world settings; however, instead of developing their own AL approach for CL, they recommend a combination of AL and CL strategies for various CL scenarios. Ayub and Fendley (2022) discuss ACL in a few-shot situation; however, they choose examples based on their proposed uncertainty measure, which focuses on learning new tasks. Perkonigg et al. (2021) address the need for ACL in a medical domain; however, because they assume that the distribution is gradually changing, their focus is on detecting the change in the distribution to determine when to label incoming images. Our study focuses on developing a new data informativeness measure that strikes a balance between preventing catastrophic forgetting and learning new knowledge.
@@ 3.2 Fisher Information-Based Active Learning
The target Fisher information matrix ${\bm{I}}({\bm{\theta}}^{*};U)$ quantifies the importance of parameters in modeling the distribution $D$ (or the unlabeled data pool $U$), while the candidate Fisher information matrix ${\bm{I}}({\bm{\theta}}^{*};S)$ quantifies the importance of parameters in modeling the candidate subset $S$. Intuitively, training a model using $S^{*}$ puts emphasis on properly estimating the essential parameter specified by the target Fisher information matrix. For the AL scenario, $\theta^{*}$ is approximated by the estimated parameter from the initially labeled data.
@@ 4 AccuACL: Proposed ACL Method
7:         $F_{M}\leftarrow{\bm{f}}({\bm{\theta}}_{t};M),F_{U}\leftarrow{\bm{f}}({\bm{\theta}}_{t};U)$triangleright Get Fisher information embedding (§$\S$ 4.3)
8:         ${\bm{F}}_{t}\leftarrow\lambda\cdot\overline{F_{M}}+(1-\lambda)\cdot\overline{F_{U}}$triangleright Get target Fisher information (§$\S$ 4.2)
9:         ${\mathbf{x}}\leftarrow\underset{i\in\{0,\ldots,|U|-1\}}{\operatorname*{arg\,max}^{(2b)}}\exp{(-\mathcal{D}_{\text{JS}}(\sigma(F_{U,i})||\sigma({\bm{F}}_{t})))}$triangleright Over-sample by distribution score (§$\S$ 4.5.1)
@@ 4.2 Accumulated Informativeness and Fisher-based ACL
where $U_{1:t}$ is the whole collection of unlabeled data from all encountered tasks $\{1,\dots,t\}$, and $b_{t}$ refers to the labeling budget for task $t$. The target Fisher information matrix ${\bm{I}}({\bm{\theta}}_{t};U_{1:t})$ represents the importance of parameters for all observed tasks. That is, the trace function in Eq. (8) represents the accumulated informativeness in Eq. (7). As a result, the optimal subset $S_{t}^{*}$ guarantees an accurate estimate of essential parameters for all observed tasks.
@@ 4.3 Fisher Information Embedding
For deep neural networks with numerous parameters, directly solving Eq. (8) is infeasible owing to the computationally expensive and time-consuming calculation of the inverse of the Fisher information matrix, which is cubic in complexity. Recent research, such as BAIT (Ash et al., 2021), reduces the computational cost by using the last linear classification layer to obtain the Fisher information matrix and an online approach to update the inverse matrix via Woodbury matrix identity. However, since matrix inversion is still needed, it is computationally demanding, confining its application to small-scale datasets such as MNIST (LeCun et al., 1998) and CIFAR-10 (Alex, 2009).
@@ 4.4 Fisher-Optimality-Preserving Properties
Property 1. Position-Wise Optimality: If we fix all Fisher embeddings $s_{i}:\forall i\neq k$, a larger $s_{k}$ will be closer to optimizing Eq. (13). That is, having a large information generally for all $s_{i}:\forall i\in[0,|{\bm{\theta}}_{t}|]$ is beneficial for optimization.
@@ 5 Experiment
5.1 Experiment Setting
Algorithms. We compare AccuACL with six AL algorithms in combination with four rehearsal-based CL methods. The implementation details can be found in Appendix G.
  • Active learning: (1) Uniform randomly selects examples from the unlabeled data at each task, (2) Entropy (Settles, 1995) selects the examples for which the model is the least certain—i.e., by selecting those with the highest entropy of predicted probability distribution, (3) LeastConfidence (Wang and Shang, 2014) selects the examples that have the smallest confidence on the highest probability class, (4) kCenterGreedy (Sener and Savarese, 2017) tries to select the most diverse examples that are farthest from each other in feature space, (5) BADGE (Ash et al., 2019) is a hybrid approach that selects both diverse and uncertain examples, (6) BAIT (Ash et al., 2021) reduces the complexity of Fisher information-based AL, making it suitable for use in deep learning environments, and Full is the fully-supervised setting.
@@ 5.1 Experiment Setting
Datasets. We use three popular CL benchmark datasets in class-incremental scenarios, SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet, which are all derived from the original CIFAR10, CIFAR100 (Alex, 2009), and TinyImageNet (Le and Yang, 2015), respectfully. SplitCIFAR10 splits 50K training images in CIFAR10 into five tasks, where each task includes 10K images for 2-way classification. SplitCIFAR-100 splits 50K training images in CIFAR100 into ten tasks, where each task includes 5K images of 10-way classification. SplitTinyImageNet divides 100K training images in TinyImageNet into ten tasks, where each task involves 10K images of 20-way classification.
@@ 5.2 Main Results
Figure 3: Effect of the parameter $\lambda$.
Table 2: Baselines with memory information.
@@
Overall Performance. Table 1 shows the overall performance of AccuACL and six AL baselines, combined with four rehearsal-based CL methods. Overall, AccuACL achieves the best performance in most cases, outperforming other AL baselines by $23.8\%$ and $17.0\%$ in average accuracy and forgetting, respectively in average. This superior performance indicates that considering the ability to prevent forgetting is as significant as the ability to acquire new knowledge quickly in query strategy to boost the performance in ACL. On the other hand, most AL baselines that focus only on learning new tasks often perform worse than Uniform. Also, BAIT is not scalable to large datasets with many classes, including SplitCIFAR100 and SplitTinyImageNet, because of its expensive query strategy. In contrast, owing to our efficient query algorithm, AccuACL succeeds in scaling to these datasets.
@@
Moreover, Figure 2(a) illustrates the performance trend of average accuracy as the per-task labeling budget varies. The results consistently show that AccuACL outperforms the AL baselines across different labeling budget settings, highlighting its robustness and adaptability. Interestingly, performance appears to reach an optimal point even without using the entire dataset for training, as seen in several settings in Table 1. This phenomenon aligns with findings in core-set selection for CL (Yoon et al., 2022), where the goal is to select high-quality subsets from a larger pool of data to enhance performance. In the context of ACL, AccuACL can be viewed as selecting high-quality examples in an unsupervised manner. The experimental settings in which ACL surpasses CL are mostly attributed to AccuACL, indicating AccuACL’s effectiveness in identifying sample-wise quality.
Efficiency.
@@
Figures 2(b) and 2(c) depict the memory usage and time spent for querying, respectively, for different AL algorithms, tested on SplitCIFAR10($M$=100). Due to our effective approximation in Section 4.3 and querying algorithm in Section 4.5, AccuACL can perform querying with reasonable efficient time and space consumption. Conversely, BAIT requires 462.1 times more time and 85.5 times more space resources than AccuACL.
5.3 Ablation Studies

### v2

@@ Abstract
Continual learning (CL) enables deep neural networks to adapt to ever-changing data distributions. In practice, there may be scenarios where annotation is costly, leading to active continual learning (ACL), which performs active learning (AL) for the CL scenarios when reducing the labeling cost by selecting the most informative subset is preferable. However, conventional AL strategies are not suitable for ACL, as they focus solely on learning the new knowledge, leading to catastrophic forgetting of previously learned tasks. Therefore, ACL requires a new AL strategy that can balance the prevention of catastrophic forgetting and the ability to quickly learn new tasks. In this paper, we propose AccuACL, Accumulated informativeness-based Active Continual Learning, by the novel use of the Fisher information matrix as a criterion for sample selection, derived from a theoretical analysis of the Fisher-optimality preservation properties within the framework of ACL, while also addressing the scalability issue of Fisher information-based AL. Extensive experiments demonstrate that AccuACL significantly outperforms AL baselines across various CL algorithms, increasing the average accuracy and forgetting by $23.8\%$ and $17.0\%$, respectively, on average.
@@ 1 Introduction
(a) Accumulated informativeness in ACL.             (b) AL performance on CL benchmarks.
Figure 1: Overview of AccuACL; (a) Unlike conventional AL strategies that only focus on the new task and cause catastrophic forgetting, AccuACL balances the prevention of catastrophic forgetting and the ability to learn new tasks quickly, by defining the accumulated informativeness; (b) shows the catastrophic forgetting of the conventional AL strategies on SplitCIFAR100.
@@
To answer the aforementioned question, we formulate the accumulated informativeness as a novel standard for informativeness in ACL, which balances the prevention of catastrophic forgetting and the ability to quickly learn new tasks. It enables assessing an example’s informativeness with respect to both past and new tasks. Then, we introduce AccuACL (Accumulated informativeness-based Active Continual Learning), an algorithm based on the theoretical analysis of the Fisher information-based AL (Sourati et al., 2017, Zhang and Oles, 2000), that maximizes the accumulated informativeness. We model the accumulated informativeness via the Fisher information matrix, through the approximation with a small memory buffer commonly adopted by many rehearsal-based CL methods, the model parameters, and the unlabeled data pool for the new task, as illustrated in Figure 1(a). As a result, AccuACL becomes to prefer the examples that comprehensively contain the four important features (i.e., Features 1$\sim$4) in Figure 1(a). Furthermore, because Fisher information-based AL is a combinatorial optimization problem, we develop a query algorithm based on two theoretical properties that an optimal labeled subset should possess. To the best of our knowledge, this is the first ACL study that offers the ability to avoid catastrophic forgetting.
Extensive experiments with four different CL methods on three CL benchmarks, SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet, show that AccuACL significantly boosts CL approaches, outperforming conventional AL baselines by $23.8\%$ and $17.0\%$, in terms of average accuracy and forgetting, respectively, on average. As shown in Figure 1(b), AccuACL always achieves dominance in both metrics throughout the entire sequence of tasks. The source code is available at https://github.com/kaist-dmlab/AccuACL.
@@ 2 Related Work
Active Learning (AL). AL is a research field that focuses on the selection of unlabeled data points for labeling by an oracle, which provides supervision, especially in domains where labeling requires significant costs (Ren et al., 2021, Tharwat and Schenck, 2023). This process allows us to optimize model performance under labeling budget constraints. Many strategies exist for selecting informative examples from unlabeled data. Uncertainty-based approaches measure model prediction uncertainty and select the most uncertain examples (Roth and Small, 2006, Settles, 1995, Wang and Shang, 2014). Diversity-based methods prioritize diverse examples to reduce redundancy between selected examples and capture a broader range of patterns and complexities in the data pool (Sener and Savarese, 2017). Hybrid approaches use gradient space embedding to select diverse and uncertain examples (Ash et al., 2019). Fisher information-based methods measure the asymptotic value of unlabeled data with theoretical guarantee (Sourati et al., 2017, Zhang and Oles, 2000). Notably, Kirsch and Gal (2022) demonstrate that modern AL algorithms optimize the same Fisher-based objective.
@@
Continual Learning (CL). CL has gained interest as a way to adapt models to new tasks over time. Numerous research has investigated different methods to address the stability-plasticity dilemma (Mermillod et al., 2013). Rehearsal-based methods store (Aljundi et al., 2019, Buzzega et al., 2020, Caccia et al., 2022, Rahaf and Lucas, 2019, Rolnick et al., 2019, Liang and Li, 2024, Kim et al., 2023a, Lin et al., 2024) or generate (Shin et al., 2017) a subset of examples from past tasks at a constant cost. These examples are then replayed for the new task to retain past knowledge. Regularization-based methods discourage significant changes to model parameters that are essential for past tasks (Aljundi et al., 2018, Kirkpatrick et al., 2017). Dynamic-structure-based methods generate distinct modules to augment the ability to learn new tasks (Yan et al., 2021, Zhou et al., 2023, Wang et al., 2023). Rehearsal-free or prompt-based methods are gaining popularity in the CL field, using pre-trained models and prompt tuning to adjust to data distribution shifts (Wang et al., 2022a; b).
@@
Active Continual Learning (ACL). There has been a lack of research effort on ACL. Vu et al. (2023) support our goal of tackling the impracticality of fully-supervised datasets in real-world settings; however, instead of developing their own AL approach for CL, they recommend a combination of AL and CL strategies for various CL scenarios. Ayub and Fendley (2022) discuss ACL in a few-shot situation; however, they choose examples based on their proposed uncertainty measure, which focuses on learning new tasks. Perkonigg et al. (2021) address the need for ACL in a medical domain; however, because they assume that the distribution is gradually changing, their focus is on detecting the change in the distribution to determine when to label incoming images.
@@ 3.2 Fisher Information-Based Active Learning
The target Fisher information matrix ${\bm{I}}({\bm{\theta}}^{*};U)$ quantifies the importance of parameters in modeling the distribution $D$ (or the unlabeled data pool $U$), while the candidate Fisher information matrix ${\bm{I}}({\bm{\theta}}^{*};S)$ quantifies the importance of parameters in modeling the candidate subset $S$. Intuitively, training a model using $S^{*}$ puts emphasis on properly estimating the essential parameter specified by the target Fisher information matrix. For the AL scenario, ${\bm{\theta}}^{*}$ is approximated by the estimated parameter from the initially labeled data.
@@ 4 AccuACL: Proposed ACL Method
7:         ${\bm{F}}_{M}\leftarrow{\bm{f}}({\bm{\theta}}_{t};M),{\bm{F}}_{U}\leftarrow{\bm{f}}({\bm{\theta}}_{t};U)$triangleright Get Fisher information embedding (§$\S$ 4.3)
8:         $F_{t}\leftarrow\lambda\cdot\overline{{\bm{F}}_{M}}+(1-\lambda)\cdot\overline{{\bm{F}}_{U}}$triangleright Get target Fisher information (§$\S$ 4.2)
9:         ${\mathbf{x}}\leftarrow\underset{i\in\{0,\ldots,|U|-1\}}{\operatorname*{arg\,max}^{(2b)}}\exp{(-\mathcal{D}_{\text{JS}}(\sigma(F_{U,i})||\sigma(fF_{t})))}$triangleright Over-sample by distribution score (§$\S$ 4.5.1)
@@ 4.2 Accumulated Informativeness and Fisher-based ACL
where $U_{1:t}$ is the whole collection of unlabeled data from all encountered tasks $\{1,\dots,t\}$, $b_{t}$ is the labeling budget for task $t$, and ${\bm{\theta}}_{t}^{*}$ is the true model parameter such that ${\bm{\theta}}_{t}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}}\mathbb{E}_{D_{1:t}}[p(y|{\mathbf{x}};{\bm{\theta}})]$. The target Fisher information matrix ${\bm{I}}({\bm{\theta}}_{t};U_{1:t})$ represents the importance of parameters for all observed tasks. That is, the trace function in Eq. (8) represents the accumulated informativeness in Eq. (7). As a result, the optimal subset $S_{t}^{*}$ guarantees an accurate estimate of essential parameters for all observed tasks.
@@ 4.3 Fisher Information Embedding
For deep neural networks with numerous parameters, directly solving Eq. (8) is infeasible owing to the computationally expensive and time-consuming calculation of the inverse of the Fisher information matrix, which is cubic in complexity. Recent research, such as BAIT (Ash et al., 2021), reduces the computational cost by using the last linear classification layer to obtain the Fisher information matrix and an online approach to update the inverse matrix via Woodbury matrix identity. However, since matrix inversion is still needed, it is computationally demanding, confining its application to small-scale datasets such as MNIST (LeCun et al., 1998) and CIFAR10 (Alex, 2009).
@@ 4.4 Fisher-Optimality-Preserving Properties
Property 1. Position-Wise Optimality: If we fix all Fisher embeddings $s_{i}:\forall i\neq k$, a larger $s_{k}$ will be closer to optimizing Eq. (13). That is, having large information generally for all $s_{i}:\forall i\in[0,|{\bm{\theta}}_{t}|]$ is beneficial for optimization.
@@ 5 Experiment
5.1 Experimental Setup
Algorithms. We compare AccuACL with six AL algorithms in combination with four rehearsal-based CL methods. The implementation details can be found in Appendix E.
  • Active learning: (1) Uniform randomly selects examples from the unlabeled data at each task, (2) Entropy (Settles, 1995) selects the examples for which the model is the least certain—i.e., by selecting those with the highest entropy of predicted probability distribution, (3) LeastConfidence (Wang and Shang, 2014) selects the examples that have the smallest confidence on the highest probability class, (4) kCenterGreedy (Sener and Savarese, 2017) tries to select the most diverse examples that are farthest from each other in feature space, (5) BADGE (Ash et al., 2019) is a hybrid approach that selects both diverse and uncertain examples, (6) BAIT (Ash et al., 2021) reduces the complexity of Fisher-based AL, making it suitable for use in deep learning environments, and Full is the fully-supervised setting.
@@ 5.1 Experimental Setup
Datasets. We use three popular CL benchmark datasets in class-incremental scenarios, SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet, which are all derived from the original CIFAR10, CIFAR100 (Alex, 2009), and TinyImageNet (Le and Yang, 2015), respectfully. SplitCIFAR10 splits 50K training images in CIFAR10 into five tasks, where each task includes 10K images for 2-way classification. SplitCIFAR100 splits 50K training images in CIFAR100 into ten tasks, where each task includes 5K images of 10-way classification. SplitTinyImageNet divides 100K training images in TinyImageNet into ten tasks, where each task involves 10K images of 20-way classification.
@@ 5.2 Main Results
Figure 3: Effect of the parameter $\lambda$ for SplitCIFAR10 on task 4.
Table 2: Performance of baseline AL methods with memory buffer examples as an initially selected subset.
@@
Overall Performance. Table 1 shows the overall performance of AccuACL and six AL baselines, combined with four rehearsal-based CL methods. Overall, AccuACL achieves the best performance in most cases, outperforming other AL baselines by $23.8\%$ and $17.0\%$ in average accuracy and forgetting, respectively in average. This superior performance indicates that considering the ability to prevent forgetting is as significant as the ability to acquire new knowledge quickly in query strategy to boost the performance in ACL. On the other hand, most AL baselines that focus only on learning new tasks actually under-perform even Uniform because they are biased too much towards new tasks, confirming our motivation. Also, BAIT is not scalable to large datasets with many classes, including SplitCIFAR100 and SplitTinyImageNet, because of its expensive query strategy. In contrast, owing to our efficient query algorithm, AccuACL succeeds in scaling to these datasets. For further analysis on the impact of task ordering, please refer to Appendix F.1.
@@
Moreover, Figure 2(a) illustrates the performance trend of average accuracy as the per-task labeling budget varies. The results consistently show that AccuACL outperforms the AL baselines across different labeling budget settings, highlighting its robustness and adaptability. Interestingly, AccuACL reaches the optimal point even without using the entire dataset for training, which is also shown in several settings in Table 1. This phenomenon aligns with findings in core-set selection for CL (Yoon et al., 2022), where the goal is to select high-quality subsets from a larger pool of data to enhance performance. In the context of ACL, AccuACL can be viewed as selecting high-quality examples in an unsupervised manner. The experimental settings in which ACL surpasses CL are mostly attributed to AccuACL, indicating AccuACL’s effectiveness in identifying example-wise quality. For further analysis on the labeling budget in a single task, please refer to Appendix F.2.
@@
Efficiency. Figures 2(b) and 2(c) depict the memory usage and time spent for querying, respectively, for different AL algorithms, tested on SplitCIFAR10($M$=100). Due to our effective approximation in Section 4.3 and querying algorithm in Section 4.5, AccuACL can perform querying with reasonable time and space consumption. Conversely, BAIT requires 462.1 times more time and 85.5 times more space resources than AccuACL.
5.3 More Results and Ablation Studies